# Disclosure of AI-Generated News Increases Engagement but Does Not Reduce Aversion, Despite Positive Quality Ratings

### v2

@@ 6 Conclusion
The results regarding our second set of outcomes are particularly noteworthy. We found significant and positive effects on respondents’ willingness to continue reading the articles after the true source was revealed, applicable to both treatment groups (AI-assisted and AI-generated). This suggests that readers may not be opposed to engaging with AI-generated or AI-assisted texts and may even be curious to learn more about AI. However, it is important to consider that the increased interest in reading the full article may not be a lasting effect changing preferences for AI-generated news. Participants might have become curious about AI involvement after learning the true authorship, but they have not become more (nor less) open towards reading AI news in general. Future research could explore this increased interest using more targeted intervention research designs, which could provide valuable insights for news producers on how to enhance specific target audiences’ interest and increase article engagement.
@@
Regarding respondents’ willingness to read AI-generated news articles in the future, we found no significant differences between the treatment groups. Simply reading two AI-assisted or AI-generated texts does not appear to alleviate concerns about AI involvement in news production, as overall willingness scores remained below 2.5 across all groups. These findings may indicate a gradual familiarization process with AI among readers. Over time and with further technological advancements, AI might become more integrated into journalism, particularly in news production, with readers becoming more accustomed to it. Therefore, future research should continue monitoring public attitudes toward AI to detect potential familiarization effects.
@@
Future research could focus on tracking changes in public perception and engagement with AI-generated news over time through longitudinal studies, which would provide insights into whether initial curiosity leads to long-term acceptance or trust. Exploring the role of transparency in disclosing AI involvement would also be important, as different types of transparency might influence how readers perceive and trust AI-generated content. Intervention studies could test different strategies to enhance reader engagement to determine what factors increase willingness to engage with AI-produced news. Expanding this research to include diverse cultural and linguistic contexts would be important, to assess whether these findings are generalizable across different populations. Additionally, studying the impact of AI-generated news on public knowledge could reveal how well readers retain and understand information from AI versus human-written articles.

### v3

@@ 6 Conclusion
The results regarding our second set of outcomes are particularly noteworthy. We found significant and positive effects on respondents’ willingness to continue reading the articles after the true source was revealed, applicable to both treatment groups (AI-assisted and AI-generated). This suggests that readers may not be opposed to engaging with AI-generated or AI-assisted texts and may even be curious to learn more about AI. This finding is in contrast with the concern that readers would disapprove of automated news, leading news organizations to refrain from disclosing that a story was automatically generated (Graefe and Bohlken, 2020), yet mirrors recent research showing that recognizing the benefits of AI in journalism, such as improved objectivity and accuracy, positively influences trust and engagement (Morosoli et al., 2024). However, it is important to consider that the increased interest in reading the full article may not be a lasting effect changing preferences for AI-generated news. Participants might have become curious about AI involvement after learning the true authorship, but they have not become more (nor less) open towards reading AI news in general. Future research could explore this increased interest using more targeted intervention research designs. This could provide valuable insights for news producers on how to enhance specific target audiences’ interest and increase article engagement, especially now that numerous news organizations are setting up guidelines to govern how they use AI (Becker, Simon and Crum, 2025).
@@
Regarding respondents’ willingness to read AI-generated news articles in the future, we found no significant differences between the treatment groups. Simply reading two AI-assisted or AI-generated texts does not appear to alleviate concerns about AI involvement in news production, as overall willingness scores remained below 2.5 across all groups. These findings may indicate a gradual familiarization process with AI among readers. Over time and with further technological advancements, AI might become more integrated into journalism, particularly in news production, with readers becoming more accustomed to it. However, our study suggests that audience hesitation is not purely a matter of unfamiliarity but is also tied to deeper normative expectations about journalistic authenticity.
Beyond its empirical findings, this study contributes to theoretical discussions on media trust, audience behavior, and the evolving role of AI in journalism. Notably, our results indicate a decoupling of content quality from audience acceptance, reinforcing the idea that trust in journalism is not solely determined by article quality but also by perceptions of authorship. This aligns with the concept of algorithmic aversion, wherein readers are reluctant to accept AI-generated content despite recognizing its quality (Dietvorst, Simmons and Massey, 2015). The study extends source credibility theory by demonstrating that while AI-generated articles may score high in expertise, they may still be viewed as less trustworthy due to the absence of human editorial oversight, accountability, and perceived authenticity (Altay and Gilardi, 2024).
These findings also contribute to ongoing debates about transparency in journalism. While transparency is often promoted as a mechanism to build trust, our results suggest that disclosing AI involvement in news production does not necessarily lead to increased long-term acceptance, highlighting the need for more refined models of how AI disclosure interacts with audience heuristics and expectations. Future research should investigate how different framing strategies—such as emphasizing human oversight or explaining AI’s editorial role—might mitigate trust concerns.
@@
Future research could focus on tracking changes in public perception and engagement with AI-generated news over time through longitudinal studies, which would provide insights into whether initial curiosity leads to long-term acceptance or trust. Exploring the role of transparency in disclosing AI involvement would also be important, as different types of transparency might influence how readers perceive and trust AI-generated content. Intervention studies could test different strategies to enhance reader engagement to determine what factors increase willingness to engage with AI-produced news. Expanding this research to include diverse cultural and linguistic contexts would be important, to assess whether these findings are generalizable across different populations. Additionally, studying the impact of AI-generated news on public knowledge could reveal how well readers retain and understand information from AI versus human-written articles. Finally, our findings emphasize the importance of balancing AI integration in journalism with audience expectations, suggesting that future news production models may benefit from hybrid approaches that maintain a human editorial presence while leveraging AI for efficiency.
Ethics Approval
This study was approved by the Ethics Committee of the Faculty of Arts and Social Sciences, University of Zurich (approval number 23.10.14).